hiiiihiiiiiiiiii:

Subtract this problem using mixed numbers. Reduce your answer to the lowest terms. 20 5/6 - 5 5/12 15 1/6 15 5/12 15 15 1/12

hiiiihiiiiiiiiii:

i thank its D

justus:

Just a second..

hiiiihiiiiiiiiii:

ok

hiiiihiiiiiiiiii:

is it D

hiiiihiiiiiiiiii:

so B

justus:

20*6+5/6 -5 5/12 120+5/6 -5 5/12 125/6 - 5 5/12 125/6 - 5*12+5/12 125/6 - 60+5/12 125/6 - 65/12 LCD is 12 125*2/6*2 -65/12 250/12 -65/12 250-65/12 185/12 15 5/12

justus:

Yes, B is the right answer.

darkknight:

A quicker way to do this is to convert 5/6 into 10/12 Now 20-5 = 15 10/12-5/12 = 5/12 Now adding them together you would get 15 5/12